# A quasi-linear model of electromagnetic turbulent transport and its application to flux-driven transport predictions for STEP

### v1

@@ 5.2 Nonlinear analysis
Figure 21: (a) Turbulent (red dots), neoclassical (blue dots) and total (green dots) power (a) and particle (b) losses at $r/a=0.5$, $r/a=0.65$ and $r/a=0.8$ of the new STEP profiles, compared to the expected heating power and particle fueling from the original STEP case (black dots). The green shaded area corresponds to a $\pm$20of the total heat (a) and particle (b) transport. The turbulent fluxes are evaluated from GENE nonlinear gyrokinetic simulations and the neoclassical fluxes are evaluated from NEO simulations, both considering the T3D steady state profiles.
At $r/a=0.8$, the quasi-linear model overestimates the heat and particle fluxes from the nonlinear simulation, despite turbulent transport being solely driven by hybrid-KBMs. Turbulent fluxes should therefore be described well by the quasi-linear model. The disagreement might be due to some inaccuracy affecting the quasi-linear model, as visible in Fig. 10 where a few points exhibit order of magnitude differences between the quasi-linear and the nonlinear heat flux predictions. We performed an additional nonlinear simulation with $a/L_{n_{e}}$, $a/L_{n_{i}}$, $a/L_{T_{e}}$, $a/L_{T_{i}}$ and $\beta^{\prime}$ increased by 5% , which returns heat and particle turbulent fluxes that are an order of magnitude larger than in the nominal case. The local equilibrium at $r/a=0.8$ lies near a marginal state where a small increase in drive can boost the turbulent fluxes significantly. Although this marginal condition (where simulations are more difficult and nonlinear effects intrinsically important) might not be described well by the quasi-linear model, transport stiffness should allow particle and power balance to be restored with a modest increase in the drive and minimal impact on the overall density and temperature profiles. In fact, transport stiffness makes flux-driven simulations less sensitive to input parameters and improves the robustness of the steady state profile predictions.
In summary, Fig. 21 shows that gyrokinetic transport fluxes for the T3D transport steady state in STEP (from T3D), are broadly consistent with the quasi-linear reduced transport model of Eqs. (8) and (9). This gives confidence in the validity of the reduced model within the parameter space explored in this T3D transport simulation. It is important to stress, however, that this is a first iteration to develop a suitable reduced model for STEP: further guidance from higher-fidelity nonlinear gyrokinetic simulations will be required to overcome important limitations of the reduced transport model presented here, including the absence of fast $\alpha$ effects. Nevertheless the reduced model and T3D-GS2 are powerful tools that could be further developed and used more extensively, together with existing integrated modelling codes, to assist in the design of new STEP plasmas and flat-top operating points.
@@ 6 Conclusions
In the reference T3D simulation, quasi-linear heat fluxes are computed approximately 300 times at each radial location before the transport steady state is reached, at a total computational cost of approximately $10^{5}$ core-hours. If fluxes were computed using nonlinear local gyrokinetics, this computational cost would have exceeded $10^{8}$ core-hours as a single calculation requires more than $10^{5}$ cores-hours in STEP-like regimes. Starting from the reduced model transport steady state from T3D as an initial condition, however, should significantly reduce the computational cost of the nonlinear gyrokinetic approach to such a flux-driven calculation. It can be envisaged that scenario design for future machines, including STEP, will benefit from exploiting a hierarchy of tools similar to that outlined in Fig. 22: integrated modelling codes like JETTO, with lower computational cost and fidelity transport models, providing initial profiles; refinement of transport predictions with intermediate fidelity physics-based models like the quasi-linear reduced model in T3D-GS2; and further honing with higher fidelity flux-driven codes like GENE-TANGO, T3D-GX or TGYRO-CGYRO, based on nonlinear gyrokinetic fluxes. This approach may also be used to improve the fidelity of models used in integrated modelling, e.g. by implementing the reduced transport model of the intermediate approach directly into JETTO (which would require either using a fast linear eigensolver or building a surrogate model).
@@ Acknowledgements
We would like to thank T. Adkins, E. Belli, J. Candy, B. Chapman-Oplopoiou, A. Di Siena, T. Görler, P. Ivanov, H. Meyer, and F. Sheffield for useful discussions. This work has been funded by the Engineering and Physical Sciences Research Council (grant numbers EP/R034737/1 and EP/W006839/1) and by STEP, a UKAEA programme to design and build a prototype fusion energy plant and a path to commercial fusion. Simulations have been performed on the ARCHER2 UK National Supercomputing Service under the project e607, on the Cambridge Service for Data Driven Discovery (CSD3) facility operated by the University of Cambridge Research Computing Service that is partially funded by the Engineering and Physical Sciences Research Council (grant EP/T022159/1), and on the Viking cluster that is a high performance compute facility provided by the University of York. To obtain further information on the data and models underlying this paper please contact PublicationsManager@ukaea.uk.

### v2

@@ 5.2 Nonlinear analysis
Figure 21: Turbulent (red dots), neoclassical (blue dots) and total (green dots) power (a) and particle (b) losses at $r/a=0.5$, $r/a=0.65$ and $r/a=0.8$ of the new STEP profiles, compared to the expected heating power and particle fueling from the original STEP case (black dots). The green shaded area corresponds to a $\pm$20of the total heat (a) and particle (b) transport. The turbulent fluxes are evaluated from GENE nonlinear gyrokinetic simulations and the neoclassical fluxes are evaluated from NEO simulations, both considering the T3D steady state profiles.
At $r/a=0.8$, the quasi-linear model overestimates the heat and particle fluxes from the nonlinear simulation, despite turbulent transport being solely driven by hybrid-KBMs. Turbulent fluxes should therefore be described well by the quasi-linear model. The disagreement might be due to some inaccuracy affecting the quasi-linear model, as visible in Fig. 10 where a few points exhibit order of magnitude differences between the quasi-linear and the nonlinear heat flux predictions. In addition, the local equilibrium at $r/a=0.8$ might be near a marginal state where a small increase in drive can boost the turbulent fluxes significantly: this marginal condition, where simulations are more difficult and nonlinear effects intrinsically important, might not be described well by the quasi-linear model, although further investigations, which are outside the aim of this work, would be required before drawing firm conclusions. We note that transport stiffness should allow particle and power balance to be restored with a modest increase in the drive and minimal impact on the overall density and temperature profiles. In fact, transport stiffness makes flux-driven simulations less sensitive to input parameters and improves the robustness of the steady state profile predictions.
In summary, Fig. 21 shows that gyrokinetic transport fluxes for the T3D transport steady state in STEP are broadly consistent with the quasi-linear reduced transport model of Eqs. (8) and (9). This gives confidence in the validity of the reduced model within the parameter space explored in this T3D transport simulation. It is important to stress, however, that this is a first iteration to develop a suitable reduced model for STEP: further guidance from higher-fidelity nonlinear gyrokinetic simulations will be required to overcome important limitations of the reduced transport model presented here, including the absence of fast $\alpha$ effects. Nevertheless the reduced model and T3D-GS2 are powerful tools that could be further developed and used more extensively, together with existing integrated modelling codes, to assist in the design of new STEP plasmas and flat-top operating points.
@@ 6 Conclusions
In the reference T3D simulation, quasi-linear heat fluxes are computed approximately 300 times at each radial location before the transport steady state is reached, at a total computational cost of approximately $10^{5}$ core-hours. If fluxes were computed using nonlinear local gyrokinetics, this computational cost would have exceeded $10^{8}$ core-hours as a single calculation requires more than $10^{5}$ cores-hours in STEP-like regimes. Starting from the reduced model transport steady state from T3D as an initial condition, however, should significantly reduce the computational cost of the nonlinear gyrokinetic approach to such a flux-driven calculation. It can be envisaged that scenario design for future machines, including STEP, will benefit from exploiting a hierarchy of tools similar to that outlined in Fig. 22: integrated modelling codes like JETTO, with lower computational cost and fidelity transport models, providing initial profiles; refinement of transport predictions with intermediate fidelity physics-based models like the quasi-linear reduced model in T3D-GS2; and further honing with higher fidelity flux-driven codes like TANGO-GENE, T3D-GX or PORTALS-CGYRO, based on nonlinear gyrokinetic fluxes. This approach may also be used to improve the fidelity of models used in integrated modelling, e.g. by implementing the reduced transport model of the intermediate approach directly into JETTO (which would require either using a fast linear eigensolver or building a surrogate model).
@@ Acknowledgements
We would like to thank T. Adkins, E. Belli, J. Candy, B. Chapman-Oplopoiou, A. Di Siena, T. Görler, P. Ivanov, H. Meyer, and F. Sheffield for useful discussions.
Funding
This work has been funded by the Engineering and Physical Sciences Research Council (grant numbers EP/R034737/1 and EP/W006839/1) and by STEP, a UKAEA programme to design and build a prototype fusion energy plant and a path to commercial fusion. Simulations have been performed on the ARCHER2 UK National Supercomputing Service under the project e607, on the Cambridge Service for Data Driven Discovery (CSD3) facility operated by the University of Cambridge Research Computing Service that is partially funded by the Engineering and Physical Sciences Research Council (grant EP/T022159/1), and on the Viking cluster that is a high performance compute facility provided by the University of York.
Declaration of interests
The authors report no conflict of interest.
Data availability
To obtain further information on the data and models underlying this paper please contact PublicationsManager@ukaea.uk.